### v1

# The role of synthetic data in Multilingual, Multi-cultural AI systems: Lessons from Indic Languages

Abstract
Developing AI systems that operate effectively across languages while remaining culturally grounded is a long-standing challenge, particularly in low-resource settings. Synthetic data provides a promising avenue, yet its effectiveness in multilingual and multicultural contexts remains underexplored. We investigate the creation and impact of synthetic, culturally contextualized datasets for Indian languages through a bottom-up generation strategy that prompts large open-source LLMs ($\geq 235$B parameters) to ground data generation in language-specific Wikipedia content. This approach complements the dominant top-down paradigm of translating synthetic datasets from high-resource languages such as English. We introduce Updesh111[Uncaptioned image]https://hf.co/datasets/microsoft/Updesh_beta, a high-quality large-scale synthetic instruction-following dataset comprising 9.5M data points across 13 Indian languages, encompassing diverse reasoning and generative tasks with an emphasis on enhancing long-context and multi-turn capabilities, in addition to improving alignment with Indian cultural contexts. A comprehensive evaluation incorporating both automated metrics and human annotation across 10k assessments indicates that the generated data is of high quality; however, human evaluation highlights specific areas for further improvement. Additionally, we perform downstream evaluations by fine-tuning models on our dataset and assessing their performance across 15 diverse multilingual datasets to evaluate the generalizability of our dataset. Our experiments demonstrate that models trained on Updesh consistently obtain significant performance improvements on generative tasks and remain competitive on multiple-choice evaluations in NLU tasks. Further, the relative improvements of models fine-tuned with Updesh are most pronounced in low and medium-resource languages, effectively narrowing the gap between these languages and high-resource languages. These findings provide empirical evidence that effective multilingual AI development requires multi-faceted data curation and generation strategies that include context-aware, culturally grounded methodologies.
1 Introduction
Building multilingual and multicultural AI is essential to ensure equitable access to technology in diverse linguistic and cultural communities. Without this inclusivity, AI systems risk reinforcing global power imbalances by privileging dominant languages and Western-centric worldviews. However, even frontier language models often underperform in languages other than English and in non-Western contexts. This is not only due to limited linguistic and cultural diversity in pre-training data, but also to English-centric practices in fine-tuning, evaluation, and other parts of the development pipeline. While pre-training data is often treated as abundant and readily available due to large-scale web crawling, fine-tuning and evaluation data must be deliberately constructed and curated to ensure quality and relevance. This often leads to skipping the creation of truly multilingual and culturally grounded data altogether, or relying on simplified approaches like translating existing English datasets, often overlooking linguistic nuance and cultural context.
As highlighted by Joshi et al. (2020), there is a significant data gap between high-resource and low-resource languages in the internet data, the major source of pre-training data. Their taxonomy classifies world languages into six resource classes, with Classes 5-6 encompassing over $2400$ languages (93.8793.87% of all world languages) serving $1.2$ billion speakers, yet having minimal to no digital resources, creating a massive representation gap in current AI systems. This gap is even more pronounced for fine-tuning and evaluation datasets (Hu et al., 2025). Synthetic data generation has become popular to augment training sets and create benchmarks, by generating new data or by transforming existing data into desirable formats, such as converting raw text into structured instruction-style prompts. Synthetic data, created mainly in English, has shown promise in improving downstream performance in tasks such as reasoning (Goldie et al., 2025; Harsha et al., 2025), coding (Wei et al., 2024; Shao et al., 2025) and retrieval (Bonifacio et al., 2022; Dai et al., 2023; Chitale et al., 2025). In this work, we study whether synthetic data has the potential to supplement scarce multilingual, multicultural data when carefully designed.
Techniques for generating synthetic data are typically grounded in English-centric practices, often relying on assumptions about model quality and reliability that do not hold for many other languages. Assessing the quality and utility of synthetic data is an ongoing area of research. Current methods for evaluating quality include human and automated evaluation to assess fluency, correctness, and diversity, which may not suffice for assessing the quality of multilingual, multicultural synthetic data. Downstream evaluation involves fine-tuning models using synthetic data and evaluating models on standardized benchmarks. In this paper, we introduce the first framework for multilingual and multicultural synthetic data generation, focusing on key dimensions such as choice of generation strategies, grounding in language-specific data, quality assessment, and evaluation. We apply this framework to a case study that involves generating synthetic data across 13 Indian languages and conduct comprehensive evaluations to check for quality and utility. Our detailed analysis across multiple multilingual synthetic datasets offers practical insights for developing more inclusive, linguistically diverse, and culturally attuned language technologies.
2 Framework
Synthetic data is utilized throughout the AI lifecycle, including pre-training, supervised fine-tuning (SFT), reinforcement learning from human feedback (RLHF), and evaluation (Viswanathan et al., 2025). Synthetic data generation starts with a small set of human-curated seeds, which is then provided to an LLM to generate instructions, inputs and outputs, followed by filtering for correctness and quality. An alternate approach is to retrieve the relevant datasets and use an LLM to rewrite them into the desired format. The dominant method for producing multilingual synthetic data involves translating English data with Neural Machine Translation systems or Large Language Models (LLMs) for pre-training (Doshi et al., 2024; Khan et al., 2024), Supervised Fine Tuning (SFT) (Shaham et al., 2024), and evaluation. However, translation may introduce translation artifacts (referred to as translationese) (Zhang & Toral, 2019; Vanmassenhove et al., 2021) and yield weaker correlations with human evaluations compared to benchmarks tailored for particular languages and contexts (Kreutzer et al., 2025; Wu et al., 2025). Recent work has explored bottom-up approaches that rely on grounding and retrieval augmentation, including instruction backtranslation (Li et al., 2024) and instruction back-and-forth translation (Nguyen et al., 2024), which generate high-quality synthetic data grounded in web corpora. These methods can potentially demonstrate superior performance compared to traditional translation-based approaches. However, there remains limited exploration of such approaches in multilingual contexts. In this work, our focus is on curating multilingual, multicultural synthetic data for SFT, although our framework can apply to other kinds of multilingual, multicultural synthetic data, including synthetic pre-training data and preference data as well. Below, we describe the key factors to consider when generating this type of data.
Base model capability: Synthetic data is usually generated by prompting large foundation models due to their fluency, ability to follow instructions, and safety guardrails. For generating multilingual, multicultural synthetic data, candidate base models should be evaluated on established multilingual and multicultural benchmarks to determine performance in target languages. If benchmarks for a specific target language are unavailable, evaluation results from related languages or overall multilingual performance can be used as indicators. Some models list supported languages and may include multilingual benchmark results in their release reports, while others do not provide these details. Additional factors to consider when selecting a base model include licensing, cost, and whether models are open-weights or restricted.
Seed data selection: Seed data is typically selected on the basis of factors such as diversity and target task coverage. Additional considerations for multilingual, multicultural seed data collection would include prioritizing tasks with coverage of cultural knowledge, norms, and values, and tasks with more relevance in specific regional and cultural contexts (for example, agriculture).
Data generation strategy: Several data generation strategies can be considered for synthetic data generation, including translation, back-translation, and retrieval-augmented synthetic data generation. Translating high-quality, curated English SFT datasets allows us to carry over critical skills the LLM needs to acquire into the multilingual training data. Back-translation using existing multilingual datasets can enhance diversity. Retrieval-augmented generation from diverse sources can potentially leverage native speaker-authored content to produce synthetic data that includes cultural knowledge and linguistic nuances.
Quality metrics: Synthetic data should be evaluated for correctness and quality before being used in downstream tasks. This may involve both automated assessment and human evaluation by native speakers. Essential quality metrics for multilingual and multicultural synthetic data include:
  • Language correctness: Is the data in the right language, register, and dialect, or is there confusion with a similar language? (Marchisio et al., 2024)
  • Linguistic acceptability: Does the data sound fluent and natural to a native speaker of the language? (Hada et al., 2024b; a)
  • Cultural appropriateness: Does the data contain appropriate references to cultural artifacts and knowledge and accurately reflect values and norms?
  • Bias and safety: Does the data amplify stereotypes, contain unsafe content, or culturally inappropriate outputs?
Downstream evaluation The primary objective of generating synthetic data is to improve model performance in real-world applications, making the choice of downstream tasks for evaluations critical. This is particularly challenging given the scarcity of multilingual and multicultural benchmarks. To address this, downstream benchmarks should cover all target languages and, wherever possible, avoid relying on translated datasets originally in English. They should also include diverse tasks, especially when synthetic data spans multiple domains. They should incorporate tests of cultural knowledge, norms, and values. In some cases, new benchmarks aligned with the target task, language, and cultural context may need to be developed to accurately assess the utility of synthetic data. Finally, care must be taken to avoid benchmark contamination in the base model, a common issue with multilingual benchmarks (Ahuja et al., 2024).
Native speaker consent, involvement and privacy: Multilingual and multicultural synthetic data creation processes should involve native speakers in seed data selection and evaluation when feasible. Informed consent is required to address cultural considerations and data sovereignty according to local regulations. Synthetic data must not contain personally identifiable information derived from seed data or online sources.
3 Related Work
English Instruction Fine-tuning Datasets: Instruction Fine-tuning (IFT) (Ouyang et al., 2022) adapts pre-trained language models to follow human instructions through supervised learning on instruction-output pairs. English IFT datasets include the FLAN collection(Wei et al., 2022), which scaled instruction tuning to over 1,800 tasks and chain-of-thought prompting, and Self-Instruct (Wang et al., 2023), which demonstrated effectiveness of LLM-generated synthetic data, leading to Stanford Alpaca (Taori et al., 2023) and the enhanced Alpaca-GPT4 (Peng et al., 2023). Subsequent work expanded through conversational approaches like Vicuna (Chiang et al., 2023) using ShareGPT conversations, WizardLM’s Xu et al. (2024) Evol-Instruct methodology, and LIMA’s (Zhou et al., 2023a) demonstration that 1K curated prompts suffice for effective tuning. The ORCA series (Mukherjee et al., 2023; Mitra et al., 2023) introduced explanation tuning and Prompt Erasure techniques, culminating in OrcaAgent-Instruct (Mitra et al., 2024a) with 25.8M synthetic instruction-response pairs.
Multilingual Instruction Fine-tuning Datasets: Multilingual instruction-following capabilities have been developed through three primary methodologies. Translation-based methods create multilingual datasets by machine translation of existing English instruction data. Bactrian-X (Li et al., 2023) translates Alpaca (Taori et al., 2023) and Dolly (Conover et al., 2023) to produce 3.4M pairs across 52 languages, mAlpaca (Chen et al., 2024) demonstrates cost-effective multilingual tuning through Alpaca translation, and LIMA-X achieves 9.9% cross-lingual improvements using $\sim$1K instructions. Template-based approaches like M2Lingual leverage Evol-guided taxonomy (Xu et al., 2024) for synthetic generation across 70 languages. Hybrid strategies combine multiple techniques: Aya-Collection (Singh et al., 2024b) integrates crowdsourcing across 65 languages with templating and translation, repurposing xP3 (Muennighoff et al., 2023), Flan (Longpre et al., 2023), and Dolly datasets while employing NLLB 3.3B (Team et al., 2022) for n-way translation. Similarly, IndicAlign aggregates 74.7M prompt-response pairs across 20 Indian languages through aggregation of existing datasets, translations via the IndicTrans2 (Gala et al., 2023) model, synthetic conversation generation from India-centric Wikipedia, and crowdsourcing.
Data Generation Strategies: Most prior efforts rely on output distillation from advanced LLMs like GPT-4, where models learn from teacher-generated responses. However, recent approaches have explored alternative strategies to address distillation limitations. Instruction Backtranslation (Li et al., 2024) starts with a model fine-tuned on seed data to generate instruction prompts for web documents, followed by self-curation to select high-quality examples. This leverages web content diversity while avoiding over-reliance on teacher models through iterative self-augmentation and self-curation. Instruction Back-and-Forth Translation (Nguyen et al., 2024) extends backtranslation by generating synthetic instructions and then rewriting responses using an LLM to improve quality based on initial documents. This combines web information diversity with high-quality response generation, outperforming distillation and yielding better win rates than ShareGPT and Alpaca-GPT4 (Peng et al., 2023).
Limitations of prior multilingual IFT Datasets: Existing multilingual instruction datasets suffer from critical limitations that hinder effective cross-lingual alignment. Translation-based approaches like Bactrian-X, mAlpaca, and LIMA-X focus primarily on basic instruction-following while excluding advanced reasoning capabilities and lack demographic or cultural representation for local alignment. These datasets suffer from translation inaccuracies due to context-length limitations of traditional sentence-level translation methods (NLLB 3.3B and IndicTrans2) which fail to capture linguistic nuances and introduce errors that can potentially creep in during training. Despite impressive language coverage, Aya-Collection provides limited culturally-specific content, while IndicAlign sources a large proportion of data from WordNet (Miller, 1994) with a focus on question-answering offering limited task diversity. Most existing datasets are short-context and single-turn, failing to address real-world usage patterns including long-context or multi-turn dialogue capabilities. This can significantly constrain model performance in practical applications that demand sustained, contextually-aware interactions.
4 Data generation
The Updesh dataset comprises two complementary subsets addressing distinct aspects of multilingual instruction following: reasoning data and open-domain generative data. This design recognizes that reasoning capabilities are largely language- and culture-agnostic, making translation-based approaches suitable for tasks like mathematical problem-solving and logical inference (Shaham et al., 2024). Existing high-quality reasoning datasets such as OrcaAgent-Instruct and OrcaMath, thus are valuable resources for multilingual adaptation. However, generative capabilities requiring cultural awareness, linguistic naturalness, and factual grounding in local contexts cannot be adequately addressed through translation due to inherent Western-centric biases and lack of cultural specificity in existing datasets (Yao et al., 2024). Therefore, our generative subset employs a grounded approach that ensures factuality through Wikipedia-based content, maintaining linguistic naturalness through native language generation, and preserves cultural adherence through systematic curation of India-specific cultural artifacts.
Reasoning Data: Inspired by prior work (Ahuja et al., 2025; Khan et al., 2024), we translate eight subsets of the OrcaAgent-Instruct (Mitra et al., 2024a) and OrcaMath (Mitra et al., 2024b) datasets into 13 Indic languages. Specifically, we consider seven reasoning-related subsets from OrcaAgent-Instruct [^2] https://huggingface.co/datasets/microsoft/orca-agentinstruct-1M-v1 along with the Math subset from OrcaMath [^3] https://huggingface.co/datasets/microsoft/orca-math-word-problems-200k (Table 1). Both datasets have been attributed to induce significant chain-of-thought and reasoning capabilities in models during instruction-tuning without the need for specific preference optimization. We employ Llama-3.1-405B-Instruct for selective translation given its strong coverage in Indian languages and instruction-following capabilities that enable adaptation to various conversational styles (Sankar et al., 2025). Decoding is performed using nucleus sampling with top_p=0.95top\_p=0.95 and temperature set to $1.0$. After generation, all outputs undergo a series of quality control filters, described in detail below.
Table 1: Reasoning task categories and their descriptions
Task Type Description
Analytical Reasoning MCQ-style questions requiring step-by-step logical inference
Multiple-Choice Questions General-purpose problems across diverse knowledge domains
Fermi (Guesstimation) Open-ended estimation problems using logical assumptions
Few-Shot Chain-of-Thought Tasks with 4-5 in-context examples for learning
Brain Teasers Puzzles stimulating lateral thinking and creativity
Text Classification Categorization tasks for predefined labels
Math Grade-school arithmetic, algebra, and geometry word problems
Open-Domain Generative Data: Synthesizing generative data presents greater challenges than translating existing content, as it introduces increased risks of hallucinations, factual inaccuracies, and demographic misalignment. We conducted a comparative analysis between Llama-3.1-405B-Instruct and Qwen3-235B-A22B models across both reasoning and non-reasoning paradigms to evaluate qualitative differences in generated outputs. Manual assessment revealed that the Qwen3-235B-A22B model demonstrated superior proficiency in generative tasks and excelled in complex instruction following owing to its strong reasoning traces. Therefore, we used the Qwen3-235B-A22B for our generative data curation process. Drawing inspiration from instruction backtranslation techniques (Li et al., 2024), our methodology begins with unlabelled content to construct questions, followed by LLM-generated answers. To ensure generation diversity and contextual grounding that maintains factual accuracy and demographic relevance, we leverage Wikipedia pages in respective target languages as our knowledge base. Table 2 summarizes the eight generative task categories. As demonstrated, some tasks required two phases, which indicates two LLM inferences in order to generate the synthetic data instance.
Table 2: Generative task categories with synthesis methods, phases, and model configuration
Task Type Synthesis Method Phases Qwen3-Mode
Logical Reasoning Generate implicit inferences from text passages (1) Direct inference generation Reasoning
Multi-Hop QA Create questions requiring information synthesis across text segments (1) Question generation (2) Answer generation Reasoning
Creative Writing Transform factual content into engaging narratives (1) Generate creative piece (2) Generate eliciting prompt Reasoning
Multi-Turn Dialogue Agentic workflows with 3-5 turn conversations between personas (1) Generate dialog adhering to personas (2) Generate natural prompt Non-reasoning
Summarization Generate summaries preserving key information across languages (1) Direct summary generation Non-reasoning
Machine Translation Cross-lingual conversion maintaining cultural context (1) Direct translation Non-reasoning
Causal Reasoning Identify and explain cause-effect relationships in text (1) Direct causal analysis Reasoning
Further, to ensure cultural representation in synthetic data generation, we systematically curated culturally relevant content from Wikipedia using the MediaWiki API. Following (Yao et al., 2024)’s cultural taxonomy framework, we traversed Wikipedia category structures starting from Category:Culture of India and Category:Culture of India by state or union territory, exploring 2-3 levels deep. This yielded two complementary datasets: a general collection (54 categories, 19,143 artifacts) covering pan-Indian elements, and a state-wise collection (28 states + 8 union territories, 37,227 artifacts) with region-specific content spanning festivals, cuisine, traditional arts, architecture, and religious practices. We sampled 26.8K cultural artifacts from this collection to create multi-hop question-answer pairs for synthetic data generation.
Data Filtering After generating approximately 7.5 million data points across 13 languages for both the Reasoning and Open-Domain subsets, manual validation was not feasible, therefore, following the approach of Shen et al. (2025), we employed automated quality checks but use the standard threshold-based method instead of their anomaly detection-based method. Specifically, we applied two filtering criteria: (1) Language Identification using IndicLID (Madhani et al., 2023) with a 0.75 confidence threshold, and (2) word repetition ratio capped at 0.75 to flag low-quality generations. Table 3 shows the filtering results, demonstrating high data quality with drop rates below 2% for most subsets. The main exception is the FS-CoT subset for Urdu, where the outputs showed excessive repetition leading to higher filtering rates, but we maintain these thresholds to ensure data integrity. For the English-to-XX translation tasks, Assamese had the highest drop rate as the model frequently generated Bengali text instead, likely due to the shared script and similarity between these languages, and because Assamese is a low-resource language.
Table 3: Document filtering statistics averaged over all Indic Languages
Reasoning Subset Generative Subset
Category Total Drop Rate (%) Category Total Drop Rate (%)
analytical_reasoning 25,000 0.047 logical_reasoning 16,384 1.459
brain_teaser 50,000 0.043 multihop_QA 16,384 1.459
fermi 25,000 0.015 creative_writing 16,384 1.459
fs_cot_flow 25,000 3.769 Multi_turn_Dialog 16,384 1.611
math 200,035 0.035 summarization 16,384 1.526
mcq 99,986 0.135 machine_translation_en 16,384 0.641
rc 50,000 0.379 machine_translation_xx 16,384 17.047
text_classification 50,000 1.878 cultural_multihop_reasoning 26,834 0.347
causal_reasoning 16,384 1.453
5 Data Quality Analysis
5.1 Translation Quality Assessment for the Reasoning Subset
For the reasoning subset, we conducted large-scale selective translation utilizing the Llama-3.1-405B-Instruct. Given that the input often consists of long-context and non-standard text, it is crucial to rigorously evaluate the quality of the generated translations. To this end, following established methodologies the generated output in the target language is translated back into English and compared with the original English source. Due to the substantial volume of data, we randomly sample 4,096 instances from each subset for every language to ensure a representative evaluation. For the backtranslation step, we employed the more efficient Llama-3.3-70B-Instruct, which offers faster inference and serves as a conservative lower bound on translation quality. We assessed translation fidelity using standard metrics such as ChrF (Popović, 2015), implemented using the SACREBLEU library (Post, 2018) for consistency. Table 9 in the Appendix demonstrates consistently high backtranslation scores across all languages and subsets, indicating robust translation quality.
5.2 Data Quality Assessment for Generative Data
Although large language models (LLMs) have emerged as scalable evaluators following the popular LLM-as-a-judge paradigm, their reliability in culturally nuanced and low-resource language settings still remains limited  (Watts et al., 2024). Our study combined LLM-based assessment with native-speaker human annotation to quantify data quality and establish inter-annotator agreement. We employed a stratified sampling strategy to select 100 data points per task category: Creative Writing, Cultural Multi-Hop Reasoning, Multi-Turn Dialogue, and aggregated reasoning tasks (Logical, Causal, and Multi-Hop Reasoning) across five Indian languages (Assamese, Gujarati, Hindi, Malayalam and Punjabi) resulting in a total of 2000 data points. Our sampling methodology ensures representative coverage by partitioning responses into quintiles based on the output length and maintaining proportional sampling across buckets to preserve the original length distribution while capturing variance across responses. Human evaluations were conducted by native speakers through an external agency (Table 7). We employed GPT-4o as an automated evaluator using identical evaluation protocols to allow direct comparison. We devised the following metrics to capture multidimensional quality aspects:
  • Creative Writing: Instruction adherence, linguistic fluency, narrative coherence
  • Reasoning Tasks: Answer adequacy, contextual adherence, instruction compliance, fluency and readability, problematic content detection, cultural relevance assessment (cultural multi-hop only)
  • Multi-Turn Dialogue: Persona consistency, topical coherence, linguistic plausibility, repetition detection, toxicity screening, instruction adherence
  • Summarization: Content coverage, factual accuracy, conciseness, coherence and logical flow, stylistic appropriateness
  • Translation: Semantic correctness, fluency corectness, domain appropriateness, style and tone, completeness
Further details on the metrics and their definitions could be found in the Appendix in Table LABEL:tab:rubrics. All metrics utilize a three-point Likert scale (0-2) with detailed rubrics provided to both human and automated evaluators to ensure assessment consistency. The prompts used for the LLM evaluations could be found in the Appendix A.7. The guidelines given to the human annotators have been added in the supplementary material. Quantitative analysis of human evaluations reveals very high data quality: across 10,000 individual metric assessments, human evaluators assigned only 27 scores of zero (0.27%), indicating that the overwhelming majority of generated content meets acceptable quality thresholds. We plan to publicly release all evaluation data, including both human and GPT assessments, to ensure complete transparency and support future research on calibration of LLM-based evaluators.
Figure 1: Human LLM-judge agreement across evaluation metrics, revealing differences across dimensions.
Inter-Annotator Agreement: To quantify the reliability of automated evaluation, we computed the percentage agreement between the LLM-judge and the human evaluators. We find significant variance in agreement across metrics (Figure 1). Agreement substantially deteriorates for culturally and linguistically nuanced assessments, particularly for linguistic plausibility evaluation and in repetition detection in long dialogue sequences. Conversely, well-defined objective criteria, such as toxicity detection and problematic content identification, demonstrate robust inter-annotator alignment. Since our data is generated using benign prompts and topics, toxic or problematic content is expected to be very rare or nonexistent. These results are consistent with earlier research on the challenges of subjective versus objective evaluation and reveal shortcomings in current LLM-judges for assessing culturally sensitive content  (Watts et al., 2024). The distribution of scores across tasks and languages for both humans and LLM can be found in Appendix  A.2.
6 Downstream Tasks Evaluation
We selected two base models, Llama-3.1-8B and Phi4-14B, for fine-tuning experiments. These models were chosen based on their size (for feasibility of experiments given available resources) and reported multilingual capabilities (Grattafiori et al., 2024; Abdin et al., 2024). We used the Axolotl framework [^4] https://github.com/axolotl-ai-cloud/axolotl for all fine-tuning runs. Details regarding the hyperparameters and compute resources used can be found in Appendix A.4.
Baselines: We fine-tuned both Llama-3.1-8B and Phi4-14B on three high-quality, open-source IFT datasets that collectively cover nearly all the Indic languages included in our work: the Aya-Collection (Singh et al., 2024b), IndicAlign (Khan et al., 2024) and the Bactrian-X dataset (Li et al., 2023). To the best of our knowledge, these are the only open datasets that offer both broad language coverage and instruction-following format, and there is a lack of other large-scale open IFT datasets for Indian languages. The Bactrian-X dataset provides training samples for 10 out of our 13 target languages, excluding Assamese, Kannada and Punjabi. The Aya-Collection dataset includes all target languages except Punjabi. Since Aya-Collection contains millions of samples per language, we uniformly sub-sampled it to construct a balanced dataset of approximately 7 million samples which is comparable in scale to Updesh. For IndicAlign, the WordNet subset ($\sim$97M pairs) is disproportionately large, less diverse, and redundant, so we downsampled it to one instance per entry, yielding about 7.3M training pairs while retaining all other subsets.
Downstream Tasks: Our evaluation framework consists of three task categories to comprehensively assess model capabilities. Natural language understanding (NLU) tasks use multiple-choice questions to measure comprehension and reasoning through likelihood-based scoring. Natural language generation (NLG) tasks like translation and summarization, assess models’ ability to generate coherent and contextually appropriate outputs. For instruction-following evaluation, we translate established benchmarks (IFEval, IFBench) using GPT-4o (prompts in A.6), enabling assessment of both cross-lingual instruction-following (English responses to non-English instructions) and native language instruction-following. This design systematically identifies model strengths and weaknesses across diverse tasks, providing a holistic performance assessment. Dataset details are in Appendix A.5.
6.1 Results
Table 4: NLU task performance. All entries are accuracy (higher is better).
Model Setting MMLU MMLU-I MILU ARC-I BoolQ-I TVQA-I Bele INCL GMMLU
Llama Base 0.23 0.23 0.26 0.25 0.62 0.27 0.24 0.25 0.23
Bactrian 0.63 0.37 0.42 0.26 0.65 0.66 0.41 0.39 0.42
Indic-Align 0.60 0.34 0.42 0.27 0.71 0.59 0.32 0.34 0.39
Aya 0.59 0.33 0.39 0.28 0.76 0.51 0.47 0.34 0.37
Updesh 0.51 0.34 0.37 0.30 0.78 0.46 0.51 0.31 0.37
Phi4 Base 0.78 0.43 0.52 0.30 0.80 0.68 0.55 0.45 0.55
Bactrian 0.78 0.40 0.48 0.29 0.74 0.67 0.48 0.42 0.48
Indic-Align 0.77 0.45 0.51 0.32 0.76 0.62 0.52 0.45 0.54
Aya 0.78 0.45 0.53 0.33 0.73 0.72 0.61 0.48 0.53
Updesh 0.78 0.49 0.54 0.32 0.81 0.60 0.74 0.47 0.57
Performance on NLU Tasks: Table 4 shows the performance of both base models fine-tuned on all three IFT datasets, along with the performance of the base models on NLU tasks. Phi4-Updesh stands out as the most consistently strong configuration, attaining the best overall scores on MMLU-I, MILU, BoolQ-I, BeleBele, INCL, and GlobalMMLU. For the rest of the tasks, different models excel - in case of the Llama base model, Bactrian-X leads knowledge-intensive tasks, attaining the highest on MMLU and TVQA-I, while Updesh excels in instruction-style reasoning benchmarks (ARC-I, BoolQ-I) and performs best on reading comprehension tasks such as BeleBele. Phi4’s MMLU performance appears saturated across all fine-tuning baselines, suggesting a ceiling effect from pre-training knowledge or lack of coverage within all the training mixtures. Overall, these results demonstrate that different IFT datasets offer distinct advantages across various NLU capabilities.
Table 5: NLG task performance. All entries are ChrF scores (higher is better).
Model Setting XSum Flores En-XX Flores XX-En IN22-Conv-Doc
Llama Base 0.16 1.00 36.91 0.45
Bactrian 0.21 23.69 46.31 14.09
Indic-Align 9.26 25.18 2.29 10.98
Aya 0.24 22.60 0.53 6.73
Updesh 17.39 32.63 46.84 21.25
Phi4 Base 11.32 23.03 51.39 22.80
Bactrian 0.31 25.93 47.75 15.45
Indic-Align 0.27 27.48 1.77 13.82
Aya 0.37 26.51 7.63 14.01
Updesh 10.21 34.58 53.29 18.74
Performance on NLG Tasks: Table 5 demonstrates that Updesh is clearly superior to the other datasets on NLG tasks. Llama-Updesh emerges as the best configuration, achieving the highest performance across all generative tasks with substantial margins over all competing datasets as well as the base model. In case of Phi4, Updesh delivers the strongest translation performance, building effectively on the already capable base model. However, summarization results are more nuanced: the base model leads on XSum with Updesh being competitive, while the other fine-tuned models perform poorly. On long-context conversational translation (IN22-Conv-Doc) outperforms all fine-tuned variants, though it slightly lags behind the Phi4 Base model. Overall, we observe that the long-context generative training data in Updesh provides significant benefits for generative tasks.
Table 6: Instruction-following capabilities. All entries are accuracy (higher is better).
Model Setting mIFEval mIFBench
Native in English Native in English
Prompt Inst. Prompt Inst. Prompt Inst. Prompt Inst.
Llama Base 0.24 0.36 0.18 0.31 0.19 0.21 0.16 0.17
Bactrian 0.17 0.27 0.14 0.25 0.17 0.18 0.11 0.13
Indic-Align 0.16 0.24 0.14 0.23 0.17 0.19 0.16 0.17
Aya 0.13 0.22 0.11 0.21 0.18 0.19 0.18 0.19
Updesh 0.20 0.30 0.18 0.28 0.16 0.19 0.14 0.15
Phi4 Base 0.22 0.32 0.19 0.30 0.14 0.16 0.13 0.14
Bactrian 0.22 0.32 0.19 0.30 0.20 0.21 0.15 0.16
Indic-Align 0.16 0.25 0.15 0.24 0.15 0.17 0.14 0.16
Aya 0.15 0.25 0.14 0.24 0.20 0.21 0.19 0.20
Updesh 0.22 0.32 0.19 0.29 0.19 0.22 0.16 0.18
Performance on Instruction Following Tasks: Table 6 presents results on the GPT-4o translated multilingual versions of IFEval and IFBench, which evaluate fine-grained constraint adherence through highly specific requirements like character positioning, word counts, and formatting specifications - constraints notably absent in any training datasets. Native refers to prompts as well as expected output in regional language, while in English uses regional language prompts but expects English responses. Prompt-level evaluation measures whether all instructions in a prompt are followed, while instruction-level evaluation measures individual instruction adherence. Updesh emphasizes general instruction following involving task completion and output formatting, making these benchmarks an effective OOD evaluation for precise constraint adherence. For Llama, fine-tuning generally degrades performance, but Updesh shows smaller drops (mIFEval: 3-4 point lesser degradation than next best - Bactrian; mIFBench: comparable or better performance than other baselines, lags Aya on in English tasks). Phi4 exhibits different dynamics - Updesh achieves positive gains on mIFBench over the Base model (+4/+2 points Native; +1/+3 points English) while maintaining base-level performance on mIFEval, outperforming other fine-tuning approaches (Aya, Indic-Align) that show mixed or negative changes. Updesh provides robust instruction-following with minimal catastrophic forgetting compared to other baselines.
Language-wise trends: Figure 2 shows a clear trend: performance improves as language resources increase for both models and task types (NLU accuracy and NLG ChrF). We illustrate this with the resource taxonomy Joshi et al. (2020), ordered from 0 (lowest-resource) to 5 (high resource). The relative improvements with Updesh are most pronounced in low/mid-resource bands, effectively narrowing the gap to high-resource languages. The base model (without fine-tuning) is the weakest, with Bactrian-X and IndicAlign often competitive in the medium resource languages while Aya-Collection is more variable. Notably, generation benefits more from data quality: NLG improvements increase more dramatically as languages become lower-resourced, most clearly for Phi4-Updesh. Across language-resource tiers, Updesh is consistently strongest, especially at higher resource levels, reaching about $\sim$0.65 for Llama (NLU), $\sim$0.76 for Phi-4 (NLU) in terms of average accuracy, and the highest ChrF for Phi-4 (NLG).
Figure 2: NLU and NLG task performance grouped by language–resource class taxonomy from Joshi et al. (2020). Updesh yields the largest relative gains in low/mid-resource languages
6.2 Discussion
Quality evaluation: Our comprehensive quality evaluation ranges from basic filters to nuanced human evaluation. Experiments with LLM-based evaluators show that while LLM-judges do well on some metrics, they struggle with subtle multilingual and multicultural contexts. For example, they often misjudge how a language is typically spoken or written, and tend to lost track of persona consistency in longer dialogues. These issues motivate the need for scalable and reliable evaluation methods that go beyond surface-level checks.
Task-type sensitivity (NLU vs.NLG): We observe variation by task type - for instance, on NLU-style tasks (usually multiple-choice), no configuration is a clear winner across all datasets; performance depends on the base model and the specific benchmark. In contrast, on NLG-style tasks that score free-form generations, models trained on Updesh consistently perform strongly. This pattern suggests that the quality control and discourse control learned during training is transferred more directly to NLG than to choice-based NLU.
Evaluation protocol differences matter: Most NLU evaluations compute correctness via likelihood comparisons over (context, option) pairs, effectively testing discrimination among short alternatives. NLG evaluations, by design, require longer generations and measure faithfulness, adequacy, and fluency under open-ended prompting. The divergence in scoring objectives partially explains why a dataset that emphasizes long-form instruction-following (like Updesh) yields larger gains on NLG than on likelihood-based NLU. We find that the Bactrian dataset tends to perform best on NLU-style benchmarks, which can be attributed to format alignment and a larger proportion of MCQ or short-from prompts.
Distribution mismatches between train and test: Updesh contains longer-context samples and lacks short-text MCQ-style items; its task composition emphasizes multi-step, generative instruction-following. Even the MCQ subset which is part of the reasoning data in Updesh consists of long-form answers to MCQs, with close to no representation of short-text or direct answers. In contrast, many NLU benchmarks, however, are MCQ-heavy and short-context. This composition shift (task format, length, and response style) likely reduces the relative NLU gains for Updesh-trained models while amplifying NLG gains, showing that models generalize best to downstream tasks with evaluation formats that resemble their training distribution.
7 Conclusion
In this work, we examined synthetic data as a potential remedy for the scarcity of multilingual and multicultural resources. Through a comprehensive framework and systematic experiments across Indian languages, we identified effective strategies for data generation, quality assessment, and downstream evaluation, beyond English-centric norms. We built Updesh, a 9.5M 13-language IFT dataset using a culturally grounded, bottom-up pipeline. Our comprehensive evaluation spanning data generation, quality assessment (human and LLM-as-judge), and downstream tasks, revealed that synthetic data can potentially bridge resource gaps. However, no single recipe dominates across tasks and models, motivating approaches tailored to diverse linguistic and cultural contexts. We will release the Updesh dataset, evaluation protocols, and detailed analyses to enable transparent, reproducible progress.
8 Ethical Considerations
Our discussion of ethical considerations is guided by the framework proposed by Bender & Friedman (2018).
Institutional Process and Oversight
The data annotation was conducted by a third-party vendor and was approved by the Institutional Review Board of our organization and by the vendor.
Data Provenance and Quality Assurance
To mitigate potential artifacts and quality issues in the synthetic data, we implemented a rigorous quality control process. This process involved both automated evaluation with GPT-4o and manual verification by human annotators. We observed a high concordance between automated and human judgments on potentially problematic content. Furthermore, given that the data was generated using state-of-the-art large language models, the baseline incidence of such content was already substantially reduced. In a manual evaluation of 500 samples, human annotators flagged only 1 sample (0.2%) on metrics pertaining to problematic content, confirming the high quality of the resulting dataset.
Annotator Demographics
Annotators were recruited through a professional external services company. All annotators assigned to a given data point were native speakers of the language represented in the data. Table 7 summarizes the annotator demographics (education, region, age distribution, and gender). Each data worker was compensated at the rate of \$2 per data point.
Category Summary
Participants 15
Qualification Post-graduation: 7
Graduation: 8
Geography Spread across 8 Indian states
Age distribution 21–30: 7
31–40: 5
41–50: 3
Gender Female: 11
Male: 4
Table 7: Participant demographics summary
Appendix A Appendix
A.1 Rubrics used for the quality evaluation of the synthetic data
Table 8: Evaluation Rubrics for AI Tasks
Task Rubric Description
Creative Writing Instruction adherence Assesses if the output strictly follows all constraints and guidelines provided in the prompt.
Fluency Evaluates the naturalness, grammatical correctness, and readability of the generated text.
Narrative coherence Checks for logical consistency in the plot, character development, and thematic elements.
Reasoning Tasks Answer adequacy Determines if the final answer is correct, complete, and directly addresses the core question.
Context adherence Measures whether the reasoning remains faithful to the provided context, avoiding external facts.
Instruction adherence Verifies that the output’s structure, format, and steps match the user’s instructions.
Fluency and readability Assesses the clarity, logical flow, and ease of understanding of the explanation.
Problematic content and cultural relevance Scrutinizes the response for harmful stereotypes and ensures it is culturally sensitive and appropriate.
Multi-turn Dialog Persona Adherence Evaluates the model’s ability to consistently maintain a specific character or role throughout the conversation.
Topic adherence Checks if the conversation remains focused on the established topic or transitions logically.
Linguistic plausibility Assesses whether the dialogue sounds natural, human-like, and contextually appropriate.
Repetitiveness Measures the degree to which the model avoids unnecessarily repeating phrases or ideas.
Toxicity check Ensures the response is free from any offensive, harmful, or inappropriate content.
Instruction adherence Verifies that the model follows meta-instructions given by the user during the dialogue.
Summarization Coverage Determines if the summary successfully captures all essential points from the source text.
Factual accuracy Checks that the summary correctly represents the information and facts from the original document.
Conciseness Evaluates whether the summary is significantly shorter than the source while retaining critical information.
Coherence and logical flow Assesses if the summary is well-structured, logically organized, and easy to follow.
Style and tone Measures how well the summary reflects the style and tone of the original text.
Translation Semantic correctness Assesses whether the meaning, intent, and nuance of the source text are accurately conveyed.
Fluency correctness Evaluates the grammatical accuracy and naturalness of the translated text in the target language.
Domain appropriateness Checks if the terminology is correct and suitable for the specific subject matter (e.g., legal, medical).
Style and tone Determines if the translation successfully captures the original author’s writing style and emotional tone.
Completeness Verifies that the entire source text has been translated without any omissions or additions.
A.2 Results from the quality evaluations of the synthetic data
Figure 3: LLM evaluations across 5 synthetically generated tasks
Figure 4: Expert human evaluations across 5 synthetically generated tasks
Figures  3 and 4 show the distribution of the scores received for the tasks we evaluated across the languages for the tasks. Expert human evaluators have consistently given a score of 2 across languages and tasks, indicating the high quality of Updesh. The disagreements between humans and the LLM is however unclear from these plots. We thereby, performed a thorough inter annotator analysis, the details of which are present in the next figures.
Figure 5: Confusion matrices showing agreement between human and LLM evaluators
Figure 6: Agreement between human and LLM evaluators per task and language respectively
Continuing the claims made from Figure 1, Figure 5 provides a clearer view of the tasks on which humans and LLMs are likely to agree or disagree. We find that the largest disagreements occur in tasks such as assessing the linguistic plausibility of a given text in a regional language. Furthermore, LLMs struggle with evaluating long-context tasks, such as evaluating whether the same persona is maintained throughout a multi-turn conversation in a regional language. There is also a notable divergence between what human evaluators consider fluent in a relatively low-resource language and what an LLM deems fluent. In contrast, we observe considerable agreement in tasks like toxicity detection and problematic content flagging. LLMs also perform reasonably well at identifying whether a text is culturally relevant, but issues arise when the evaluation requires more fine-grained judgments of multilinguality and multiculturalism. We do not identify any specific trends language or task-wise as apparent in Figure6.
A.3 Data Quality Assesment of Reasoning Data
Language analytical brain fermi fs (cot_flow) math mcq rc text_class
Assamese 75.02 71.62 79.93 96.59 79.87 64.33 65.07 71.29
Bengali 87.25 77.40 80.69 82.10 79.87 67.25 74.94 74.72
Gujarati 77.86 67.15 82.14 73.96 49.66 78.04 63.21 55.95
Hindi 84.49 79.23 81.54 87.11 64.25 74.80 73.71 67.00
Kannada 79.96 76.87 80.02 81.26 65.80 69.91 64.15 60.77
Malayalam 75.21 73.41 70.10 77.93 68.51 55.69 63.19 75.63
Marathi 77.84 68.63 69.33 82.81 68.56 64.48 56.49 60.58
Nepali 81.79 86.18 74.71 83.32 53.98 56.42 53.86 59.63
Odia 56.47 62.06 50.70 93.61 57.55 51.21 42.95 52.15
Punjabi 83.75 51.79 77.40 79.04 61.19 51.21 69.81 54.11
Tamil 79.28 70.70 74.54 75.16 57.83 60.66 63.94 49.65
Telugu 78.24 80.26 74.33 79.91 69.88 60.66 61.93 60.95
Urdu 85.05 79.97 66.31 81.93 59.80 61.76 64.91 67.23
Table 9: Backtranslation ChrF scores for the Reasoning subset.
A.4 Hyperparameters
Table 10: Training hyperparameters used for all our experiments
Hyperparameter Value
Base Model phi4-base-hf / llama-3.1-8b-hf
Sequence Length 65,536
Effective Batch Size 8192
Number of Epochs 3
Optimizer AdamW Torch Fused
Learning Rate 1.0e-05
LR Scheduler Cosine
Adam Beta1 0.9
Adam Beta2 0.95
Max Grad Norm 1.0
Warmup Ratio 0.03
Weight Decay 0.1
NEFTune Noise Alpha 5
Precision BF16
Flash Attention True
Gradient Checkpointing True
A.5 Evaluation Datasets
Table 11: Evaluation datasets
Dataset Source
NLU Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2021)
MMLU Indic (MMLU-I) SarvamAI
ARC Indic (ARC-I) SarvamAI
BoolQ Indic (BoolQ-I) SarvamAI
TriviaQA Indic (TVQA-I) SarvamAI
BeleBele (Bele) Bandarkar et al. (2024)
INCLUDE (INCL) Romanou et al. (2025)
Global MMLU (GMMLU) Singh et al. (2025)
NLG Extreme Summarization (Xsum) Singh et al. (2024a)
Flores English to Other (Flores EnXX) Goyal et al. (2022)
Flores XX-En (Flores XXEn)
IN22-Conv (in22-conv-doc) Gala et al. (2023); Gumma et al. (2025)
IF Multilingual Instruction Following Evaluation (mIFEval) Zhou et al. (2023b)
Multilingual Instruction Following Benchmark (mIFBench) Pyatkin et al. (2025)
A.6 Prompt used for conversion of IFEval and IFBench
Translation Generation Prompt INSTRUCTION: You are a professional translator. Your sole task is to translate the provided English text into {target_language}. Rules: Output only the translation—nothing else (no explanations, no commentary). Do not answer or solve any task contained in the text; just translate it. Preserve formatting, punctuation, numbers, placeholders, and bracketed tokens exactly as they appear. Translate idiomatically and naturally for a fluent native speaker of {target_language}. If the input is an instruction (e.g., “Write a resume for …”), translate the instruction itself—do not perform it. Do not introduce extra formatting (e.g., don’t create JSON blocks unless the source already has them). SOURCE LANGUAGE: English TARGET LANGUAGE:
{target_language}
TEXT TO TRANSLATE:
{source_text}
OUTPUT REQUIREMENT:
Return only the translated text in {target_language} with formatting faithfully preserved.
A.7 Sample Prompts Used for Evaluation
The following are representative prompts (one per task) used to evaluate different task types. Each prompt follows the established rubrics from Table LABEL:tab:rubrics.
A.7.1 Creative Writing - Instruction Adherence
Creative Writing Instruction Adherence Evaluation Prompt INSTRUCTION: You are an expert literary critic and evaluator, tasked with assessing the degree to which a synthetically generated creative piece adheres to the user’s writing prompt. You are required to read the USER QUESTION thoroughly and analyze how well the generated response incorporates the specified narrative elements, stylistic choices, and constraints. USER QUESTION: {user_prompt} TARGET LANGUAGE: {tgt_lang} ASSISTANT GENERATED RESPONSE: {assistant_output} A response that adheres to the user’s creative brief has: Narrative & Thematic Completeness: It fully incorporates all requested characters, plot points, settings, and themes. The creative piece feels complete and resolves according to the prompt’s guidelines. Stylistic & Tonal Adherence: The content’s tone, mood, and writing style (e.g., genre conventions, a specific author’s voice) directly match the user’s request. Format & Constraint Compliance: It follows all explicit formatting requirements (e.g., poem, script, short story) and abides by any constraints (e.g., word count, inclusion/exclusion of specific words, use of certain literary devices). Creative Intent Alignment: It successfully captures the spirit and intended artistic goal of the user’s prompt, creating a piece that feels like a faithful realization of the user’s idea. Use the following scoring scale: 5 – Excellent: The response masterfully incorporates all creative constraints, including plot, character, tone, style, and format. It not only follows the instructions to the letter but also demonstrates a creative flair that enhances the user’s original idea. 4 – Good: The response successfully incorporates most creative instructions. There may be minor deviations in tone or style, or a secondary plot/character element might be slightly underdeveloped, but the core creative vision is clearly and effectively realized. 3 – Fair: The response addresses some of the key creative instructions but neglects or misinterprets others. For instance, it might follow the plot but fail to capture the requested tone, or it might ignore a crucial character trait or constraint. 2 – Poor: The response shows significant deviation from the creative brief. It may follow a single, simple instruction (like the general topic) but disregards crucial constraints like genre, character personality, plot structure, or mood. 1 – Unacceptable: The response completely disregards the creative instructions. The generated text is thematically, structurally, and stylistically unrelated to the user’s prompt. Return your evaluation in the following JSON format:
  "score": <integer from 1 to 5>,
  "reason": "<brief explanation for the score>"
Do not include markdown, comments, or anything outside the JSON.
A.7.2 Multi-turn Dialog - Persona Adherence
Multi-turn Dialog Evaluation Prompt INSTRUCTION: You are an expert evaluator tasked with assessing how well a multi-turn dialog maintains persona consistency. You must analyze the conversation to determine if the assistant consistently embodies the specified character or role throughout the interaction. USER QUESTION: {user_prompt} TARGET LANGUAGE: {tgt_lang} ASSISTANT GENERATED RESPONSE: {assistant_output} A response that demonstrates strong persona adherence has: Character Consistency: The assistant maintains the same personality traits, speaking style, and behavioral patterns throughout the conversation. Role-Appropriate Knowledge: The responses reflect knowledge and expertise appropriate to the specified persona. Consistent Voice: The tone, vocabulary, and manner of speaking remain true to the character across all turns. Believable Interactions: The persona feels authentic and natural in the conversational context. Use the following scoring scale: 5 – Excellent: Perfect persona consistency with natural, believable character embodiment throughout all turns. 4 – Good: Strong persona adherence with minor inconsistencies that don’t break character immersion. 3 – Fair: Generally maintains persona but has noticeable lapses or inconsistencies in character. 2 – Poor: Significant persona inconsistencies that frequently break character immersion. 1 – Unacceptable: Complete failure to maintain persona or embody the specified character. Return your evaluation in the following JSON format:
  "score": <integer from 1 to 5>,
  "reason": "<brief explanation for the score>"
Do not include markdown, comments, or anything outside the JSON.
A.7.3 Reasoning Tasks - Answer Adequacy
Reasoning Task Evaluation Prompt INSTRUCTION: You are an expert evaluator specializing in logical reasoning and problem-solving. Your task is to assess whether the generated response provides a correct, complete, and well-reasoned answer to the given question. USER QUESTION: {user_prompt} TARGET LANGUAGE: {tgt_lang} ASSISTANT GENERATED RESPONSE: {assistant_output} A response with excellent answer adequacy demonstrates: Correctness: The final answer is factually accurate and logically sound. Completeness: All aspects of the question are addressed without omitting important elements. Direct Relevance: The response directly answers what was asked without unnecessary tangents. Clear Reasoning: The logical steps leading to the conclusion are evident and valid. Use the following scoring scale: 5 – Excellent: Completely correct and comprehensive answer with clear, logical reasoning. 4 – Good: Correct answer with minor gaps in completeness or explanation clarity. 3 – Fair: Generally correct but missing some important aspects or contains minor errors. 2 – Poor: Partially correct but has significant errors or omissions in reasoning or conclusion. 1 – Unacceptable: Incorrect answer or completely fails to address the question asked. Return your evaluation in the following JSON format:
  "score": <integer from 1 to 5>,
  "reason": "<brief explanation for the score>"
Do not include markdown, comments, or anything outside the JSON.